# On the Ability of Deep Networks to Learn Symmetries from Data: A Neural Kernel Theory

### v1

@@ Abstract
Symmetries (transformations by group actions) are present in many datasets, and leveraging them holds significant promise for improving predictions in machine learning. In this work, we aim to understand when and how deep networks can learn symmetries from data. We focus on a supervised classification paradigm where data symmetries are only partially observed during training: some classes include all transformations of a cyclic group, while others include only a subset. This setup mimics real-world scenarios, such as learning object pose invariance from limited examples. We ask: can deep networks generalize symmetry invariance to the partially sampled classes?
In the infinite-width limit, where kernel analogies apply, we derive a neural kernel theory of symmetry learning to address this question. The group-cyclic nature of the dataset allows us to analyze the spectrum of neural kernels in the Fourier domain; here, we find a simple characterization of the generalization error as a function of the interaction between class separation (signal) and class-orbit density (noise). We observe that generalization can only be successful when the local structure of the data prevails over its non-local, symmetric, structure, in the kernel space defined by the architecture. This occurs when (1) classes are sufficiently distinct and (2) class orbits are sufficiently dense.
Our framework also applies to equivariant architectures (e.g., CNNs), and recovers their success in the special case where the architecture matches the inherent symmetry of the data (due to orbit collapse in kernel space). Empirically, our theory reproduces the generalization failure of finite-width networks (MLP, CNN, ViT) trained on partially observed versions of rotated-MNIST. We conclude that conventional networks trained with supervision lack a mechanism to learn symmetries that have not been explicitly embedded in their architecture a priori. In the future, our framework could be extended to guide the design of architectures and training procedures able to learn symmetries from data.
@@
Date: December 16, 2024
@@ 1 Introduction
The ability to make accurate predictions depends on how well one understands the structure of a problem. Physics, notably, has achieved remarkable success in predicting natural phenomena by capturing their structure in compact mathematical equations. A key aspect of this structure lies in the concept of symmetries (Noether,, 1918; Gross,, 1996), which can be defined as transformations by group actions that do not change the identity of objects. For example, a chair remains a chair whether it is presented upright or upside down (a transformation in SO(3)). In the context of deep learning, which is nowadays widely used for prediction tasks—such as identifying objects in images—a crucial question is whether deep networks have mechanisms to recognize and exploit data symmetries for effective predictions. For example, can deep networks learn to predict the identity of objects independently of their viewpoint? And importantly, do they need to be exposed to all object classes in all possible poses during training, or is exposure to some classes in some poses sufficient for them to capture the concept of pose invariance?
@@
The field of geometric deep learning develops theories and methods that enable neural networks to take advantage of problem symmetries (Bronstein et al.,, 2021). In particular, equivariant neural networks (Cohen and Welling,, 2016) can ensure representation equivariance and classification invariance to prespecified symmetries. However, equivariant architectures require one to know the symmetry of the problem in advance. Here, we aim to understand whether conventional deep networks, which have not been explicitly designed to capture a prespecified symmetry, can learn symmetries directly from data.
@@
We focus on a supervised classification paradigm where the symmetries of the data are only partially observed during training : for some classes, all possible transformations of a cyclic group are observed during training, while for other classes only a subset of transformations is observed. This scenario mimics realistic real-world learning problems. For example, a child sees during their development a few objects in all possible 3D poses (e.g., the toys they can manipulate), and many objects in only some poses (e.g., heavy furniture). We study whether a deep network trained on such a data diet should be expected to generalize to the partially sampled classes (e.g., recognize a piece of furniture seen from an unusual viewpoint). Specifically, we ask: under what condition(s) will the network correctly extrapolate the symmetry invariance to the partially sampled classes?
@@
To characterize the generalization capabilities of deep networks in the presence of data symmetries, we rely on the equivalence between infinitely wide neural networks and kernel machines (Neal, 1996b, ; Neal, 1996a, ; Lee et al.,, 2018; Jacot et al.,, 2018). Typical neural kernels (MLP, CNN) greatly simplify when computed over a dataset generated by a cyclic group action (they become circulant), allowing an interpretable analysis in the Fourier domain of when and how symmetries are correctly learned. We find that the generalization behavior of networks is predicted by a simple ratio of inverse kernel frequency powers computed over orbits of the cyclic group. Our analysis of this formula makes clear that deep networks (as described by their kernel equivalent) are a priori unable to leverage data symmetries for generalization. Successful generalization is yet possible, in cases where the local structure of the data prevails over symmetry considerations. This happens in particular (1) when classes are sufficiently well separated in kernel space, and (2) when the symmetric structure of the data is sufficiently local in kernel space. Finally, while there is no guaranteed equivalence between finite-width networks and their infinite-width counterpart, our spectral kernel theory captures well the behavior of trained networks on symmetric datasets in all our experiments.
@@
We first briefly review prior theoretical work and practical methods developed for deep learning in relation to symmetries (Section 2). We then illustrate a simple symmetry learning problem, where deep networks with common architectures (MLP, CNN, ViT) are trained and evaluated on partial views of rotated-MNIST, a version of the well-known handwritten character dataset (Lecun et al.,, 1998) augmented with rotations (Section 3). We then analyze theoretically how kernel machines in general—and neural networks in particular—behave on datasets presenting symmetries. We build our theory through scenarios of increasing complexity, from a simple Gaussian kernel applied to a circular dataset, to deep neural kernels with or without equivariant architectures (MLPs and CNNs) applied to rotated-MNIST. In all cases, we show both theoretically and empirically the inadequacy of conventional neural architectures trained with supervision to learn symmetries that have not been embedded in their kernel design a priori. Our work provides theoretical tools which could prove helpful in identifying and discovering architectures and training procedures able to learn symmetries from data (as discussed in Section 4).
2 Prior work on deep learning, symmetries, and kernel machines
@@
The interplay between deep learning and symmetries has been studied extensively, both empirically and theoretically, for finite-width networks and in the infinite-width limit (where kernel analogies apply). Here we attempt a condensed review of these efforts. In the following, our operational definition of dataset symmetries is a set of transformations by group actions that do not change the class identity of objects present in a dataset. These group transformations may act directly in the native space of the dataset (e.g., image translations) or in a latent space affecting the dataset (e.g., images of 3D-rotating objects).
@@
Empirically, the inability of conventional neural networks to capture symmetries present in datasets has been observed in many different contexts. Conventional networks have been shown to fail to extrapolate a simple periodic function (Ziyin et al.,, 2020). In vision, studies have investigated the generalization capabilities of deep networks to recognize objects undergoing changes in pose (Alcorn et al.,, 2019; Abbas and Deny,, 2023; Madan et al.,, 2022; Siddiqui et al.,, 2023), size (Ibrahim et al., 2022b, ), mirror symmetry (Sundaram et al.,, 2021), lighting conditions (Madan et al.,, 2021), and shown a substantial degradation of network performance in these conditions. Ollikka et al., (2024) showed in a recent study that humans beat state-of-the-art deep networks and most vision-language models at recognizing objects in unusual poses. In language, compositional generalization has also been framed as capturing permutation symmetries, and traditional language models have been shown to fail to capture this symmetry (Gordon et al.,, 2019). In our work, we study from a theoretical point of view why networks fail to generalize on data symmetries despite being exposed to them partially during training, by characterizing network behavior on symmetric datasets in the kernel limit.
@@
Network architectures have been designed to be equivariant to prespecified symmetries (Gens and Domingos,, 2014; Cohen et al.,, 2018, 2019; Worrall et al.,, 2017), or conserve certain physical quantities (Finzi et al.,, 2020; Cranmer et al.,, 2020; Greydanus et al.,, 2019). Network architectures have also been designed to respect the topology of a problem (for a review: Hajij et al.,, 2023). Relaxed and adaptive equivariance schemes have also been proposed (Elsayed et al.,, 2020; d’Ascoli et al.,, 2021; Wang et al.,, 2022; Zhou et al.,, 2021; Yeh et al.,, 2022; Kaba and Ravanbakhsh,, 2023). The failure of traditional convolutional architectures to be fully equivariant has been observed (e.g., Azulay and Weiss,, 2019), and some measure of approximate equivariance proposed (Gruver et al.,, 2023). Others have investigated how equivariant representations affect the capacity of group-invariant linear readouts (Farrell et al.,, 2021). Beyond equivariance, symmetries inherent to the architecture of deep networks have also been studied, and shown to affect their learning dynamics and solutions (Tanaka and Kunin,, 2021; Simsek et al.,, 2021; Ainsworth et al.,, 2022).
@@
The effects of training networks on symmetry-augmented versions of a dataset have also been scrutinized. The efficiency of training networks on augmented data vs. enforcing equivariance in the architecture has been carefully characterized at scale (Brehmer et al.,, 2024). Moskalev et al., (2023) clarify that training on augmented datasets does not produce genuine equivariance, in the sense that the trained networks may not be equivariant outside the training distribution. Recent theoretical results show, however, that training on a perfectly augmented dataset produces emergent equivariant representations in ensembles of networks, in and outside the training set, both in the finite (Nordenfors and Flinth,, 2024) and infinite-width limit (Gerken and Kessel,, 2024). In our work we ask a related but different question: we study the generalization capabilities of networks trained on a dataset containing unknown symmetries—where no systematic augmentation scheme or equivariant architecture has been enforced. The symmetries in the training dataset may be well represented for some classes, but not others. The question we ask is then: how well can networks generalize the symmetry invariance to these partially sampled classes?
@@
Beyond equivariant and augmentation methods to deal with prespecified symmetries, approaches to learning the symmetries present in data have been proposed (Zhou et al.,, 2021; Connor and Rozell,, 2020; Keller and Welling, 2021a, ; Sanborn et al.,, 2023; Connor et al.,, 2024; Culpepper and Olshausen,, 2009; Sohl-Dickstein et al.,, 2010; Dupont et al.,, 2020; Jaderberg et al.,, 2015) and studied theoretically (Anselmi et al.,, 2019, 2023; Pfau et al.,, 2020). Self-supervised learning approaches build invariance (Ibrahim et al., 2022a, ; Zemel and Hinton,, 1990; Chen et al.,, 2020; Zbontar et al.,, 2021) or equivariance (Garrido et al.,, 2024) to predefined symmetries by feeding augmentations of the dataset to two identical versions of the network and matching their latent representations. Disentanglement has also been framed as the problem of learning symmetries from data (Higgins et al.,, 2018; Mercatali et al.,, 2022). Subsequent work has shown, however, that topological defects result from attempting to learn disentangled representations of even the simplest symmetries such as affine transformations (Bouchacourt et al.,, 2021).
@@
Beyond symmetries, the interplay between machine learning and data geometry has been studied for many types of data structure: compositional (Liang et al.,, 2024; Schott et al.,, 2021; Wiedemer et al.,, 2024; Sabour et al.,, 2017; Lippl and Stachenfeld,, 2024), hierarchical (Saxe et al.,, 2019; Mel and Ganguli,, 2021), data lying on a manifold (Goldt et al.,, 2020; Gerace et al.,, 2022) or on separate manifolds (Sorscher et al.,, 2022; Cohen et al.,, 2020; Chung et al.,, 2018).
@@
Two main kernel theories of deep learning exist, which identify deep networks with kernel machines (Neal, 1996b, ; Neal, 1996a, ; Jacot et al.,, 2018) [^1] We refer the reader unfamiliar with neural kernel theories to the excellent lecture notes of Pr. Adityanarayanan Radhakrishnan available online on this topic (https://aditradha.com/lecture-notes/).. Both theories apply in the limit where the networks have infinite width (infinite number of neurons within a layer or within a channel for convolutional networks), and both identify training a network with doing simple Gaussian Process regression over a fixed kernel determined by the network architecture. The Neural Network Gaussian Process (NNGP) theory (Neal, 1996b, ; Neal, 1996a, ) characterizes the distribution of functions produced by random draws of the network parameters (its weights and biases), assuming a distribution from which the parameters are drawn (typically i.i.d. Normal). Conditioned on a training set, the NNGP characterizes the distribution of solutions (i.e., functions passing through the training input-label pairs) under the assumed distribution of parameters, and can be used to make predictions on a testing set. The Neural Tangent Kernel theory (NTK) (Jacot et al.,, 2018) describes the distribution of solutions obtained by training all layers with gradient descent, from random initial conditions. It is allegedly the most realistic theory (although see Avidan et al., (2023) for a more nuanced account) and the one we use in this work. Nonetheless, we repeated our analyses using the NNGP kernel instead of the NTK and it did not change our conclusions (not shown). Neural kernel theories assume networks to have infinite width, which is a big approximation in practice. For example, Fort et al., (2020) show that the kernel of finite-width networks changes rapidly during the first few epochs of training to a more favourable kernel for the task, breaking the frozen kernel assumption which is only provable in the infinite-width limit. However, in all our experiments, we find that kernel theories adequately capture the generalization behavior of finite-width trained networks on datasets presenting symmetries.
@@
Finally, in an insightful line of work, Bordelon et al., (2020); Canatar et al., (2021) relate generalization properties of infinite-width deep networks to spectral properties of their kernel. They do not however explore the specific interplay between spectral kernel properties and data symmetries. In our work, we find that the symmetric nature of the dataset considerably simplifies the spectral description of neural kernels, allowing better interpretability of the factors on which generalization depends.
3 Learning from partially observed symmetries
@@
We start with a simple empirical study which illustrates the problem of symmetry learning and highlights the failure of a range of common architectures on this problem.

### v2

# On the Ability of Deep Networks to Learn Symmetries from Data – A Neural Kernel Theory

@@ Abstract
Symmetries (transformations by group actions) are present in many datasets, and leveraging them holds considerable promise for improving predictions in machine learning. In this work, we aim to understand when and how deep networks—with standard architectures trained in a standard, supervised way—learn symmetries from data. Inspired by real-world scenarios, we study a classification paradigm where data symmetries are only partially observed during training: some classes include all transformations of a cyclic group, while others—only a subset. We ask: under which conditions will deep networks correctly classify the partially sampled classes?
In the infinite-width limit, where neural networks behave like kernel machines, we derive a neural kernel theory of symmetry learning. The group-cyclic nature of the dataset allows us to analyze the Gram matrix of neural kernels in the Fourier domain; here we find a simple characterization of the generalization error as a function of class separation (signal) and class-orbit density (noise). This characterization reveals that generalization can only be successful when the local structure of the data prevails over its non-local, symmetry-induced structure, in the kernel space defined by the architecture. This occurs when (1) classes are sufficiently distinct and (2) class orbits are sufficiently dense.
We extend our theoretical treatment to any finite group, including non-abelian groups. Our framework also applies to equivariant architectures (e.g., CNNs), and recovers their success in the special case where the architecture matches the inherent symmetry of the data. Empirically, our theory reproduces the generalization failure of finite-width networks (MLP, CNN, ViT) trained on partially observed versions of rotated-MNIST. We conclude that conventional deep networks lack a mechanism to learn symmetries that have not been explicitly embedded in their architecture a priori. In the future, our framework could be extended to guide the design of architectures and training procedures able to learn symmetries from data.
@@
Keywords: deep learning, symmetry, neural kernel methods, gaussian processes, spectral methods
@@ 1 Introduction
The ability to make accurate predictions depends on how well one understands the structure of a problem. Physics, in particular, has achieved remarkable success in predicting natural phenomena by capturing their structure in compact mathematical equations. A key aspect of this structure lies in the concept of symmetry (Noether, 1918; Gross, 1996). Symmetries describe transformations by group actions that do not affect the identity of objects. For example, a chair remains a chair whether it is presented upright or upside down (a transformation in $SO(3)$). In the context of deep learning, nowadays widely used for prediction tasks, a crucial question is whether deep networks have mechanisms to recognize and exploit data symmetries for effective predictions. For example, can deep networks learn to predict the identity of objects independently of their viewpoint? And importantly, do they need to be exposed to all object classes in all possible poses during training, or is exposure to some classes in some poses sufficient for them to capture the concept of pose invariance?
@@
The field of geometric deep learning develops theories and methods that enable neural networks to take advantage of problem symmetries (Bronstein et al., 2021). In particular, equivariant neural networks (Cohen and Welling, 2016) can ensure representation equivariance and classification invariance to prespecified symmetries. However, equivariant architectures require one to know the symmetry of the problem in advance. Here, we aim to understand whether conventional deep networks—which have not been explicitly designed to capture a prespecified symmetry—can learn symmetries directly from data.
@@
We focus on a supervised classification paradigm where the symmetries of the data are only partially observed during training : for some classes, all possible transformations of a cyclic group are observed during training, while for other classes only a subset of transformations is observed. This scenario replicates realistic real-world learning problems. For example, a child during their development sees a few objects in all possible 3D poses (e.g., the toys they can manipulate), and many objects in only some poses (e.g., heavy furniture). Should we expect a deep network trained on such a data diet to generalize to the partially sampled classes (e.g., recognize a piece of furniture seen from an unusual viewpoint at test time)? In this work, we study the conditions under which deep networks correctly extrapolate the symmetry invariance to the partially sampled classes.
@@
To characterize the generalization capabilities of deep networks in the presence of data symmetries, we rely on the equivalence between infinitely wide neural networks and kernel machines (Neal, 1996; Lee et al., 2018; Jacot et al., 2018). Typical neural kernels (MLP, CNN) greatly simplify when computed over a dataset generated by a cyclic group action (they become circulant), allowing an interpretable analysis in the Fourier domain of when and how symmetries are correctly learned. We find that the generalization behavior of networks is predicted by a simple ratio of inverse kernel frequency powers computed over orbits of the cyclic group. Our analysis of this formula makes clear that deep networks (as described by their kernel equivalent) are a priori unable to leverage data symmetries for generalization. Successful generalization is yet possible, in cases where the local structure of the data prevails over its non-local, symmetry-induced structure. This happens in particular (1) when classes are sufficiently well separated in kernel space, and (2) when the symmetric structure of the data is sufficiently local in kernel space. Importantly, while there is no guaranteed equivalence between finite-width networks and their infinite-width counterpart, our spectral kernel theory captures well the behavior of normally trained finite-width networks in all our experiments on rotated-MNIST, a version of the well-known handwritten character dataset (LeCun et al., 1998) augmented with rotations.
@@
In Section 2, we briefly review prior theoretical work and practical methods developed for deep learning in relation to symmetries. In Section 3, we illustrate a simple symmetry learning problem, where deep networks with common architectures (MLP, CNN, ViT) are trained and evaluated on partial views of rotated-MNIST. In Section 4, we analyze theoretically how kernel machines in general—and neural networks in particular—behave on datasets presenting symmetries. We build our theory through scenarios of increasing complexity, from a simple Gaussian kernel applied to a circular dataset, to deep neural kernels with or without equivariant architectures (MLPs and CNNs) applied to an affine group transformation such as rotated-MNIST. In all cases, we show both theoretically and empirically the inadequacy of conventional neural architectures trained with supervision to learn symmetries that have not been embedded in their kernel design a priori. In Section 5, we discuss how our work provides theoretical tools which could be helpful in identifying and discovering architectures and training procedures able to learn symmetries from data. In Appendix C, we extend our theory to any finite group, including non-abelian groups.
2 Prior Work on Deep Learning, Symmetries, and Neural Kernels
@@
The interplay between deep learning and symmetries has been studied extensively, both empirically and theoretically, for finite-width networks and in the infinite-width limit (where kernel analogies apply). Here we attempt a condensed review of these efforts. In the following, our operational definition of dataset symmetries is a set of transformations by group action that do not change the class identity of objects present in a dataset. These group transformations may act directly in the native space of the dataset (e.g., image translations) or in a latent space affecting the dataset (e.g., images of 3D-rotating objects).
@@
Empirically, the inability of conventional neural networks to capture symmetries present in datasets has been observed in many different contexts. Conventional networks have been shown to fail to extrapolate a simple periodic function (Ziyin et al., 2020). In vision, studies have investigated the generalization capabilities of deep networks to recognize objects undergoing changes in pose (Alcorn et al., 2019; Madan et al., 2022; Abbas and Deny, 2023; Siddiqui et al., 2023), size (Ibrahim et al., 2023), mirror symmetry (Sundaram et al., 2022), lighting conditions (Madan et al., 2025), and shown a substantial degradation of network performance in these conditions. Recently, Ollikka et al. (2025) also showed that humans beat state-of-the-art deep networks and most vision-language models at recognizing objects in unusual poses. In language, compositional generalization has also been framed as capturing permutation symmetries, which traditional language models have been shown to fail at. In our work, we study from a theoretical point of view why networks fail to generalize on data symmetries despite being exposed to them partially during training. To do so, we characterize network behavior on symmetric datasets in the kernel limit.
@@
Network architectures have been designed to be equivariant to prespecified symmetries (Gens and Domingos, 2014; Cohen and Welling, 2016; Worrall et al., 2017; Cohen et al., 2019a, b; Bekkers, 2021; Weiler et al., 2021), or conserve certain physical quantities (Greydanus et al., 2019; Cranmer et al., 2020; Finzi et al., 2020; van der Ouderaa et al., 2024). Network architectures have also been designed to respect the topology of a problem (for a review, see Hajij et al., 2023). Relaxed and adaptive equivariance schemes have also been proposed (Elsayed et al., 2020; Zhou et al., 2021; D’Ascoli et al., 2021; Wang et al., 2022; Yeh et al., 2022; Kaba and Ravanbakhsh, 2023; van der Ouderaa et al., 2023). The failure of traditional convolutional architectures to be fully equivariant has been observed (e.g., Azulay and Weiss, 2019), and some measure of approximate equivariance proposed (Gruver et al., 2023). Others have investigated how equivariant representations affect the capacity of group-invariant linear readouts (Farrell et al., 2022). Beyond equivariance, symmetries inherent to the architecture of deep networks have also been studied, and shown to affect their learning dynamics and solutions (Tanaka and Kunin, 2021; Simsek et al., 2021; Ainsworth et al., 2023). In our work, we study the ability of neural networks to generalize on datasets presenting symmetries to which they have not been designed to be equivariant.
@@
The effects of training networks on symmetry-augmented versions of a dataset have also been scrutinized. Recently, the efficiency of training networks on augmented data vs. enforcing equivariance in the architecture has been carefully characterized at scale (Brehmer et al., 2024) and on a variety of datasets (Vadgama et al., 2025). Moskalev et al. (2023) clarify that training on augmented datasets does not produce genuine equivariance, in the sense that the trained networks may not be equivariant outside the training distribution. Recent theoretical results show, however, that training on a perfectly augmented dataset produces emergent equivariant representations in ensembles of networks, in and outside the training set, both in the finite (Nordenfors and Flinth, 2024) and infinite-width limit (Gerken and Kessel, 2024). In our work we ask a related but different question: we study the generalization capabilities of networks trained on a dataset containing unknown symmetries—where no systematic augmentation scheme or equivariant architecture has been enforced. The symmetries in the training dataset may be well represented for some classes, but not others. The question we ask is then: how well will networks generalize the symmetry invariance to the partially sampled classes?
@@
Beyond equivariant and augmentation methods to deal with prespecified symmetries, approaches to learning the symmetries present in data have been proposed (Culpepper and Olshausen, 2009; Jaderberg et al., 2015; Sohl-Dickstein et al., 2017; Dupont et al., 2020; Benton et al., 2020; Connor and Rozell, 2020; Zhou et al., 2021; Keller and Welling, 2021b; Pérez Rey et al., 2023; Sanborn et al., 2023; Connor et al., 2024; van der Linden et al., 2024; Yang et al., 2024b) and studied theoretically (Anselmi et al., 2019; Pfau et al., 2020; Anselmi et al., 2023; Marchetti et al., 2024). Self-supervised learning approaches build invariance (Zemel and Hinton, 1990; Chen et al., 2020; Zbontar et al., 2021; Ibrahim et al., 2022) or equivariance (Garrido et al., 2024) to predefined symmetries by feeding augmentations of the dataset to two identical versions of the network and matching their latent representations. Disentanglement has also been framed as the problem of learning symmetries from data (Higgins et al., 2018; Mercatali et al., 2022). Subsequent work has shown, however, that topological defects result from attempting to learn disentangled representations of even the simplest symmetries such as affine transformations (Bouchacourt et al., 2021; Esmaeili et al., 2023). In our work, we focus on characterizing the symmetry learning ability of conventional architectures (MLP, CNN) trained with traditional supervised learning. However, our theoretical framework could potentially be extended to characterize the ability of these other methods to capture the symmetries of data.
@@
Beyond symmetries, the interplay between machine learning and data geometry has been studied for many types of data structure: compositional (Sabour et al., 2017; Schott et al., 2022; Liang et al., 2024; Wiedemer et al., 2023; Lippl and Stachenfeld, 2025; Yang et al., 2025), hierarchical (Saxe et al., 2019; Mel and Ganguli, 2021), data lying on a manifold (Goldt et al., 2020; Gerace et al., 2022) or on separate manifolds (Chung et al., 2018; Cohen et al., 2020; Sorscher et al., 2022). Our work focuses on datasets presenting symmetries (i.e., group-action induced structures), an important type of structure found in the physical world and present in many datasets.
@@
Two main kernel theories of deep learning exist, which identify deep networks with kernel machines (Neal, 1996; Jacot et al., 2018). [^1] We refer the reader unfamiliar with neural kernel theories to the excellent lecture notes of Adityanarayanan Radhakrishnan available online on this topic (https://aditradha.com/lecture-notes/). We also provide an intuitive explanation of neural kernel theories in Appendix A. Both theories apply in the limit where the networks have infinite width (infinite number of neurons within a layer or infinite number of channels in convolutional networks), and both identify training a network with doing simple Gaussian Process regression over a fixed kernel determined by the network architecture. The Neural Network Gaussian Process (NNGP) theory (Neal, 1996) characterizes the distribution of functions produced by random draws of the network parameters (its weights and biases), assuming a distribution from which the parameters are drawn (typically i.i.d. Normal). Conditioned on a training set, the NNGP characterizes the distribution of solutions (i.e., functions passing through the training input-label pairs) under the assumed distribution of parameters, and can be used to make predictions on a testing set. The Neural Tangent Kernel theory (NTK) (Jacot et al., 2018) describes the distribution of solutions obtained by training all layers with gradient descent, from random initial conditions. It is allegedly the most realistic theory (although see Avidan et al. (2025) for a more nuanced account) and the one we use in this work. Nonetheless, we repeated our analyses using the NNGP kernel instead of the NTK and it did not change our conclusions (not shown). Neural kernel theories assume networks to have infinite width, which is a big approximation in practice. For example, Fort et al. (2020) show that the kernel of finite-width networks changes rapidly during the first few epochs of training to a more favourable kernel for the task, breaking the frozen kernel assumption which is only provable in the infinite-width limit. However, in all our experiments, we find that kernel theories adequately capture the generalization behavior of finite-width, normally trained networks on datasets presenting symmetries.
@@
Finally, in an insightful line of work, Bordelon et al. (2020); Canatar et al. (2021) relate generalization properties of infinite-width deep networks to spectral properties of their kernel. They do not, however, explore the specific interplay between spectral kernel properties and data symmetries. In our work, we find that the symmetric nature of the dataset considerably simplifies the spectral description of neural kernels, allowing better interpretability of the factors on which generalization depends.
@@
We next propose a simple empirical study to illustrate the problem of symmetry learning and highlight the failure of a range of common architectures on this problem.
3 Illustration of a Symmetry Learning Problem